OpenStudy (anonymous):

if a/b=7/2, then 2a=?

OpenStudy (anonymous):

14?

OpenStudy (anonymous):

cause aren't you multiplying 7*2

OpenStudy (anonymous):

there's an answer for 14b? so that may be it

OpenStudy (welshfella):

cross multiply 2 *a = 7*b

OpenStudy (anonymous):

see i did the whole cross multiplying part but my answer always came out different but I may just be overlooking something simple

OpenStudy (anonymous):

is the 14b right?

OpenStudy (welshfella):

no

OpenStudy (welshfella):

2 * a = 2a 7*b = 7b

OpenStudy (welshfella):

2a = 7b

OpenStudy (anonymous):

ooooohhhh dang i'm dumb, I understand now i was dividing 7/2 then trying to multiply like a dummy